### v1

# Polynomial Inequalities and Optimal Stability of Numerical Integrators

Abstract
A numerical integrator for $\dot{x}=f(x)$ is called stable if, when applied to the 1D Dahlquist test equation $\dot{x}=\lambda x,\lambda\in\mathbb{C}$ with fixed timestep $h>0$, the numerical solution remains bounded as the number of steps tends to infinity. It is well known that no explicit integrator may remain stable beyond certain limits in $\lambda$. Furthermore, these stability limits are only tight for certain specific integrators (different in each case), which may then be called ‘optimally stable’. Such optimal stability results are typically proven using sophisticated techniques from complex analysis, leading to rather abstruse proofs. In this article, we pursue an alternative approach, exploiting connections with the Bernstein and Markov brothers inequalities for polynomials. This simplifies the proofs greatly and offers a framework which unifies the diverse results that have been obtained.
1 Introduction
For a very wide class of explicit numerical integrators for ordinary differential equations (ODEs) of the form $\dot{x}=f(x)$, optimal stability limits for the Dahlquist test equation $\dot{x}=\lambda x$ have been proven in [8]. These limits, for an integrator using $m$ evaluations of $f(x)$ per step, correspond to the three cases of interest: 1) $|1+h\lambda/m|\leq 1$ for arbitrary $\lambda\in\mathbb{C},\mathrm{Re}\{\lambda\}\leq 0$; 2) $-2m^{2}\leq h\lambda\leq 0$ for $\lambda\in\mathbb{R}_{-}$; and 3) $-im\leq h\lambda\leq im$ for $\lambda\in i\mathbb{R}$. Proving such general results requires the use of complex analysis, which unfortunately encumbers the exposition.
However, if one considers a slightly less general family of integrators (still including all explicit Runge-Kutta methods), the stability analysis may be conducted via the analysis of polynomials. In this context, it has long been common to find the corresponding ‘optimal polynomials’ simply stated as such, without proof of their optimality [1, 9, 12]. Such terseness also proves frustrating for researchers seeking a more pedagogical approach to this important special case.
The present article also restricts attention to integrators which may be analysed via considering polynomials, and we may thus avoid using complex analysis. We then detail extremely simple, expository proofs which elucidate connections between the ‘optimal polynomials’ and the Bernstein [10] and Markov Brothers’ inequalities [6, pp. 89–91]. We are thus able to collect and unify results from multiple sources for the three cases of interest, depending on $\lambda$, within the same framework, offering a resource of pedagogical value which is lacking in the current literature.
Our proofs, while less general, still apply to a great many integrators of practical interest, including all (explicit) Runge-Kutta methods. This is demonstrated by the fact that all theorems in this work offer simplified proofs of published results: theorem 1 proves the result of [7] for explicit Runge-Kutta methods; theorem 2 (appropriately adapted) proves the result of [5]; and theorem 4 unifies and clarifies the results of [14, 9, 12] and [13, Sec. 4.3].
Numerical Integration
For the problem $f:D\subset\mathbb{R}^{d}\to\mathbb{R}^{d}$,
$\frac{dx}{dt}\equiv\dot{x}=f(x),\quad x(0)=x_{0},\quad t\in[0,T],$ (1)
with exact solution flow $\phi_{t}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ one may design numerical integrators which are discrete maps that furnish approximations $\{x_{n}\}_{n=1}^{N}$ to the exact solution $x(t)$, $x_{n}\approx x(nh)$ at grid points $t_{n}=nh,n=1,\ldots N$ such that $t_{N}=T=Nh$. One such class of integrators are so-called (explicit) ‘one-step maps’ (including all explicit Runge-Kutta methods) of the form
$x_{n+1}=\psi_{h}(x_{n}),\quad x_{0}=x(0),$ (2)
which generate the approximations at the grid points in an iterative fashion. Such an integrator is called consistent if in the limit $h\to 0$ $\|\psi_{h}(x)-\phi_{h}(x)\|=\mathcal{O}(h^{p+1}),\forall x\in\mathbb{R}^{d}$ with $p\geq 1$, which is a minimal condition typically required of any numerical integrator to be considered a ‘good’ approximator. Consistency however, gives little clue as to the behaviour of the integrator for moderate timesteps, which may be more relevant in certain applications. The most important aspect of this behaviour is stability. Numerical instability refers to related undesirable phenomena which manifests when the integrator is operated with stepsizes $h$ which are too large: magnification of the rounding errors that occur in finite arithmetic; inordinate divergence from the exact solution; or failure to match the qualitative behaviour of the continuous dynamics eq. 1 near fixed points. To attempt to understand these related issues, one considers the behaviour of the integrator via the aforementioned Dahlquist test problem.
Justification of the Dahlquist test problem
We consider the relevance of the test problem to the one of the aforementioned phenomena: the magnification of rounding errors. Denote the numerical solution implemented in exact arithmetic, $x_{n}$ and the computer implementation (with roundoff errors) $\widetilde{x}_{n}$. Then one has that the error $\widetilde{E}_{n}\equiv x_{n}-\widetilde{x}_{n}$ due to roundoff evolved under the exact flow follows
$\frac{d}{dt}\widetilde{E}_{n}=f(x_{n})-f(\widetilde{x}_{n})\approx\mathrm{\textbf{D}}(x_{n}-\widetilde{x}_{n})=\mathrm{\textbf{D}}\widetilde{E}_{n},\quad\mathrm{\textbf{D}}\equiv\left.\frac{df}{dx}\right|_{x_{n}},$ (3)
where the approximation is ‘good’ for $x_{n}$ close to $\widetilde{x}_{n}$ (i.e. small roundoff error, as one would expect). Consequently one has stability with respect to perturbations due to rounding errors for the exact flow if the eigenvalues of D all have non-positive real part. It is then desirable if the evolution under the numerical map $\psi_{h}$ applied to the linearised flow eq. 3 respects this, i.e. $\|\psi_{h}(\widetilde{E}_{n})\|\leq\|\widetilde{E}_{n}\|$. If D is diagonalisable, eq. 3 reduces to $d$ copies of $\dot{E}=\lambda E,\lambda\in\mathbb{C}$ which is none other than the now standard test for (A-)stability (due to Dahlquist)
$\dot{x}=\lambda x,\quad x(0)=1,\qquad\{\lambda\in\mathbb{C}:\mathrm{Re}\{\lambda\}\leq 0\}.$ (4)
Since diagonalisation commutes with almost all methods of practical interest, one may thus study stability via eq. 4 without loss of generality.
Setup
We consider consistent, explicit numerical integrators $\psi_{h}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ applied with constant timestep $h$, which, when applied to eq. 4 give an iterative solution of the form
$x_{n+1}=P(h\lambda)x_{n},$ (5)
where $P(z)$ is a polynomial of degree $\leq m$ in $z$ with $P(z)=1+z+\mathcal{O}(z^{2})$. This is the case for e.g. any consistent $m$-stage explicit Runge-Kutta (RK) method. In general a method using $m$ evaluations of $f$ (such as an $m$-stage RK method) gives a polynomial of degree $m$. Since a method using $m$ evaluations of $f$ has the same work as a method using 1 evaluation applied $m$ times with stepsize $h/m$, in the following we always compare methods of equivalent cost $m$ (i.e. polynomials of degree $m$) [8].
It is then apparent from eq. 5 that one avoids exponential instability (i.e exponential growth of the error $x_{n}$) if $\lvert P(h\lambda)\rvert\leq 1$. One then calls the stability domain of the method with stability polynomial $P(z)$ the region
$\mathcal{S}=\{z\in\mathbb{C}:|P(z)|\leq 1\}.$ (6)
For example, the Explicit Euler method for eq. 1, $x_{n+1}=x_{n}+hf(x_{n})$, applied to the test problem eq. 4 gives $x_{n+1}=(1+h\lambda)x_{n},$ and so the stability polynomial of (Explicit) Euler is $P_{Euler}(z)=1+z$, with stability domain the complex disc $\mathcal{D}_{1}=\{z\in\mathbb{C}:|1+z|\leq 1\}.$
Note that, on the other hand, the Implicit Euler method $x_{n+1}=x_{n}+hf(x_{n+1})$, applied to the test problem eq. 4 gives $x_{n+1}=(1-h\lambda)^{-1}x_{n},$ which is stable $\forall\lambda\in\mathbb{C},\mathrm{Re}\{\lambda\}\leq 0$. It is in fact the case that unlimited stability of this sort can only be possessed by implicit methods. However, implicit methods are more costly than explicit methods, and may be unsuitable for various reasons (e.g. high dimension), so that the study of the (limited) stability of explicit methods is eminently worthwhile. Given that their stability is limited, it is of interest to ask what are the “optimal stability limits” of explicit integrators which give rise to eq. 5, and which numerical integrators attain these limits. These limits also depend on the problem at hand, which induces certain restrictions on $\lambda$:
  • for general problems, $\lambda\in\mathbb{C},\mathrm{Re}\{\lambda\}\leq 0$, the optimal stability region is $\mathcal{D}_{m}=\{\lambda h\in\mathbb{C}:|1+\lambda h/m|\leq 1\}$(theorem 1);
  • for ‘parabolic problems’ $\lambda\in\mathbb{R},\lambda\leq 0$, the optimal stability region is $-2m^{2}\leq\lambda h\leq 0$ (theorem 2);
  • for ‘hyperbolic problems’ $\lambda\in i\mathbb{R}$, the optimal stability region is $-i(m-1)\leq\lambda h\leq i(m-1)$ (theorem 4).
The names parabolic and hyperbolic come from the partial differential equations (PDE) context, in which upon discretising in space a parabolic (hyperbolic) PDE, one arrives at an equation of the form eq. 1 with Jacobian having negative (pure imaginary) spectrum [13, Sec. 1.4]. Note also that in the general and parabolic cases, the optimal limits are not enhanced by enlarging the family of integrators beyond those which give rise to eq. 5; only in the case of hyperbolic problems does one see improvement, to $-im\leq\lambda h\leq im$ (for a multistep method) [8].
2 General Theorem
Two of the most prominent families of general numerical integrators for eq. 1 are Runge-Kutta methods (which are ‘one-step’ in the sense of eq. 2) and (linear) multi-step methods. We explicitly do not consider multi-step methods, since they do not give rise to maps of the form eq. 5 when applied to eq. 4. However, the result below in theorem 1, that applies to all Runge-Kutta methods is already quite strong, and was first shown in [7], although in a less direct way.
Theorem 1 (Optimal Stability - General).
Consider a consistent, explicit integrator for eq. 1 which, when applied to eq. 4 generates a map of the form of eq. 5. Then its stability domain $\mathcal{S}_{m}$ contains the complex disc
$\mathcal{D}_{m}=\{z\in\mathbb{C}:|1+z/m|\leq 1\},$
iff $P(z)=(1+z/m)^{m}$. This may be obtained by simply applying $m$ Explicit Euler steps in succession with stepsize $h/m$, i.e. $P_{mEuler}(z)=(1+z/m)^{m}$.
Proof.
Since the stability polynomial $P(z)=1+z+\mathcal{O}(z^{2})$ of degree $\leq m$ in $z$, one has
$P(0)=P^{\prime}(0)=1.$ (7)
Consider now the bijective transformation $\mu=1+z/m$, which maps $\mathcal{D}_{m}$ to $\mathcal{D}_{1}$, so that $P(z)=\widetilde{P}(\mu)$. Since $\widetilde{P}$ is a polynomial, it is analytic (i.e. complex differentiable everywhere), as is the transformation $\mu$, and so we can differentiate and apply the chain rule safely, giving
$\frac{d}{dz}P(z)=\frac{d}{dz}\widetilde{P}(\mu)=\widetilde{P}^{\prime}(\mu)\frac{d\mu}{dz}=\widetilde{P}^{\prime}(\mu)/m.$
Since $P^{\prime}(z=0)=1$, this implies that $\widetilde{P}^{\prime}(\mu=1)=m$, which holds for any stability polynomial. We may then apply the Bernstein inequality [10] (suggested on [8, p.76])
$\max_{|\mu|=1}|\widetilde{P}^{\prime}(\mu)|\leq m\max_{|\mu|=1}|\widetilde{P}(\mu)|,$ (Bernstein)
and insert the requirement that $|\widetilde{P}(\mu)|\leq 1$ for stability, and since we know that for $\mu=1$ the derivative attains the value $m$, $\max_{\lvert\mu\rvert=1}|\widetilde{P}^{\prime}(\mu)|\geq m$, giving
$m\leq\max_{|\mu|=1}|\widetilde{P}^{\prime}(\mu)|\leq m\max_{|\mu|=1}|\widetilde{P}(\mu)|\leq m,$
and so one has equality $\max_{|\mu|=1}|\widetilde{P}^{\prime}(\mu)|=m\max_{|\mu|=1}|\widetilde{P}(\mu)|$, which holds only in the case of the Bernstein polynomial $\widetilde{P}(\mu)=\mu^{m}$, i.e. $P(z)=(1+z/m)^{m}$ [10, Sec. 2]. ∎
Remark 1.
The optimal stability of the Euler method, shown by theorem 1, does in fact hold even when one allows competitor multi-step methods and other numerical integrators that do not give rise to eq. 5, although the proof requires intricate complex analysis [8, Sec. 3].
3 Parabolic Theorem
For ‘parabolic problems’, where D in eq. 3 has real spectrum, one is interested in eq. 4 with $\lambda\leq 0,\lambda\in\mathbb{R}$. In this case, the domain of stability of relevance is the line segment $-a\leq\lambda h\leq 0$. We show now that the maximal $a=2m^{2}$ for integrators, such as Runge-Kutta methods, giving rise to eq. 5, via essentially the same simple proof as that of theorem 1.
Theorem 2 (Optimal Stability - Parabolic).
Consider a consistent, explicit integrator for eq. 1 which, when applied to eq. 4 with $\lambda\leq 0$ generates a map of the form of eq. 5. Then it attains the optimal stability interval $-2m^{2}\leq z<0$ iff $P(z)=T_{m}(1+z/m^{2})$, where $T_{m}(x)$ is the $m^{\text{th}}$ Chebyshev polynomial of the first kind.
The optimal polynomial may be obtained via a sequence of $m$ Euler steps of different stepsizes such that $P(z)=\prod_{i=1}^{m}(1+z/\xi_{i})$, where $\xi_{i}=m^{2}\left[1-\cos\left(\frac{\pi(2i-1)}{2m}\right)\right]$ [1].
Proof.
Let $z=\lambda h$ as before, and consider now the transformation $\mu=1+z/m^{2}$, which maps $[-2m^{2},0]$ to $[-1,1]$, so that $P(z)=\widetilde{P}(\mu)$. Differentiating $\widetilde{P}$ as before and using $P^{\prime}(z=0)=1$ from consistency implies that $\widetilde{P}^{\prime}(\mu=1)=m^{2}$. We may then apply the Markov Brothers inequality [6, p.91] (suggested on [8, p.79])
$\max_{-1\leq\mu\leq 1}|\widetilde{P}^{\prime}(\mu)|\leq m^{2}\max_{-1\leq\mu\leq 1}|\widetilde{P}(\mu)|,$ (Markov Brothers)
and insert the requirement that $|\widetilde{P}(\mu)|\leq 1$ for stability, and since we know that for $\mu=1$ the derivative attains the value $m^{2}$, $\max_{-1\leq\mu\leq 1}|\widetilde{P}^{\prime}(\mu)|\geq m^{2}$. Continuing as before, one has equality $\max_{-1\leq\mu\leq 1}|\widetilde{P}^{\prime}(\mu)|=m^{2}\max_{-1\leq\mu\leq 1}|\widetilde{P}(\mu)|$, which holds only in the case of the Chebyshev polynomial $\widetilde{P}(\mu)=T_{m}(\mu)$, i.e. $P(z)=T_{m}(1+z/m^{2})$ [11]. ∎
Remark 2.
The optimal stability $-2m^{2}\leq z\leq 0$ for $T_{m}(1+z/m^{2})$ holds even when allowing for a broader class of competitor integrators, including linear multi-step methods, not only those giving rise to eq. 5. As for theorem 1, this extension of the result of theorem 2 requires the use of elaborate complex analysis [8, Sec. 4].
Remark 3.
In fact theorem 2 also proves an optimality result for integrators for the second-order ODE $\ddot{x}=f(x)$, with test problem $\ddot{x}=-\omega^{2}x,\omega\in\mathbb{R}$. Our result thus recovers the result of [5], of optimal stability interval $h\omega\leq 2m$, in a simpler way. In this context, the candidate methods are not Runge-Kutta but ‘splitting’ methods, whose stability may be treated by a polynomial of degree $m$ in $z=(h\omega)^{2}$, $P(z)=1-z/2+\mathcal{O}(z^{2})$ [2]. From theorem 2, with transformation $\mu=1-z/(2m^{2})$ the optimal polynomial is then $P(z)=T_{m}(1-z/(2m^{2}))$ [3, Secs. 4.4-4.5]. Note that for the special case $\ddot{x}=-Ax+g(x)$, to which one can apply more exotic (splitting) integrators, the Chebyshev polynomials are also optimal, in an appropriate sense [4].
4 Hyperbolic Theorem
For ‘hyperbolic problems’, where D in eq. 3 has pure imaginary spectrum, one is interested in eq. 4 with $\lambda\in i\mathbb{R}$. In this case, the domain of stability of relevance is the line segment $-ia\leq\lambda h\leq ia$. We show now that the optimal $a=(m-1)$ for integrators, such as Runge-Kutta methods, giving rise to eq. 5.
Before proceeding, we first require a simple lemma.
Lemma 3.
Consider a polynomial $P(z)=1+z+\alpha z^{2}+\mathcal{O}(z^{3})$ of degree $m$ in $z$, for some unknown $\alpha\in\mathbb{R}$, such that $|P(i\sqrt{y})|\leq 1,y\in[0,a],a>0$. Then $\alpha\geq 1/2$.
Proof.
Consider $P(i\sqrt{y}),y\in\mathbb{R}$ so that
$Q(y)=P(i\sqrt{y})P(-i\sqrt{y})=\left|P(i\sqrt{y})\right|^{2}=1+(1-2\alpha)y+\mathcal{O}(y^{2}),$
is a real polynomial of degree $m$ in $y$. Necessarily $\alpha\geq 1/2$ as otherwise $Q(y)>1$ (and consequently $|P(i\sqrt{y})|>1$) for all positive $y$ less than some value sufficiently close to 0. ∎
Theorem 4 (Optimal Stability - Hyperbolic).
Consider a consistent, explicit integrator for eq. 1 which, when applied to eq. 4 with $\lambda\in i\mathbb{R}$ generates a map of the form of eq. 5. Then it attains the optimal stability interval $-i(m-1)\leq z\leq i(m-1)$ iff
$P(z)=i^{m-1}T_{m-1}\left(\frac{z}{i(m-1)}\right)+i^{m-2}\left(1+\frac{z^{2}}{(m-1)^{2}}\right)U_{m-2}\left(\frac{z}{i(m-1)}\right).$ (8)
Proof.
Let $z=\lambda h$ as before, and consider $P(z)=1+z+\alpha z^{2}+\mathcal{O}(z^{3})$, for some unknown $\alpha\in\mathbb{R}$. One then has $P(i\sqrt{y}),y\in\mathbb{R}$ such that
$P(i\sqrt{y})=R+i\sqrt{y}I,\quad R(y)=1-\alpha y+\mathcal{O}(y^{2}),\quad I(y)=1+\mathcal{O}(y),$
where for $m=2k+1$ ($m=2k$), $k\in\mathbb{N}_{0}$, $R$ is of degree $k$ ($k$) in $y$ and $I$ is of degree $k$ ($k-1$). Defining $Q(y)=|P(i\sqrt{y})|^{2}=R^{2}+yI^{2}$ as in lemma 3, it is clear that the maximal interval $0\leq y\leq a^{2}$ over which $Q(y)\leq 1$ is equal to or contained by those of $R(y)^{2}$ and $yI(y)$ separately. We consider the cases of even $m=2k$ and odd $m=2k+1$ separately.
Case $m=2k+1$
In this case we first optimise the interval $R(y)$ and then seek $\sqrt{y}I(y)$ such that the interval is preserved. Using that $R(0)=1,R^{\prime}(0)=-\alpha$ and the transformation $\mu=1-2y/a^{2}$ (which maps $0\leq y\leq a^{2}$ to $-1\leq\mu\leq 1$) to give $\widetilde{R}(\mu)=R(y)$, from theorem 2 one sees that the optimal polynomial is of the Chebyshev form and fulfills $\frac{\alpha a^{2}}{2}=k^{2},$ and thus, since from lemma 3 one has $\alpha\geq 1/2$, one has the maximal $a^{2}=4k^{2}$ obtained for $R(y)=T_{k}\left(1-y/(2k^{2})\right)$.
If we wish that $Q(y)\leq 1$, we require then that $\sqrt{y}I(y)=0$ when $R(y)^{2}=1$ (i.e. at the extrema $\xi_{j}=\cos(j\pi/k),j=0,1,\ldots,k$ of $T_{k}$), which is obtained by a polynomial of the form
$\sqrt{y}I(y)\propto\sqrt{y}\left(1-\frac{y}{4k^{2}}\right)\frac{d}{dy}T_{k}\left(1-\frac{y}{2k^{2}}\right),$
where one may use the fact that $T^{\prime}_{k}(x)=kU_{k-1}(x)$, where $U_{k}$ are the Chebyshev polynomials of the second kind, and the condition $I(0)=1$ to determine the constant of proportionality, to give the optimal polynomial
$\sqrt{y}I(y)=\frac{\sqrt{y}}{k}\left(1-\frac{y}{4k^{2}}\right)U_{k-1}\left(1-\frac{y}{2k^{2}}\right),$
where $I(y)$ is indeed of degree $k$ in $y$. Taking $\sqrt{y}/2k\equiv\cos(\theta)$, one may then verify that $Q(y)=\cos^{2}(2k\theta)+\sin^{2}(\theta)\sin^{2}(2k\theta)\leq 1,0\leq y\leq 4k^{2}$ as required, using the properties of the Chebyshev polynomials $T_{k}(\cos(\theta))=\cos(k\theta),U_{k-1}(\cos(\theta))\sin(\theta)=\cos(k\theta)$.
Consequently, one may obtain stability for $-(m-1)i\leq z\leq(m-1)i$ where $m=2k+1$ for the polynomial (replacing $z=i\sqrt{y}$) [13, p. 257]
$P(z)=T_{k}\left(1+\frac{z^{2}}{2k^{2}}\right)+\frac{z}{k}\left(1+\frac{z^{2}}{4k^{2}}\right)U_{k-1}\left(1+\frac{z^{2}}{2k^{2}}\right).$
Note that by the same substitution $\sqrt{y}/2k\equiv\cos(\theta)$, one may express this polynomial as [9]
$P(z)=(-1)^{k}\left[T_{2k}\left(\frac{z}{2ki}\right)-i\left(1+\frac{z^{2}}{4k^{2}}\right)U_{2k-1}\left(\frac{z}{2ki}\right)\right].$ (9)
Case $m=2k$
When $m$ is even one cannot proceed as before to design $I(y)$ since $I(y)$ may only be of degree $k-1$. Hence, we instead begin by optimising $\sqrt{y}I$, which is of degree $12k-$ 1 in $\sqrt{y}$, and then seek $R(y)$ such that the interval is preserved. We first apply the transformation $\mu=\sqrt{y}/(2k-1)$ (which maps $-(2k-1)\leq\sqrt{y}\leq(2k-1)$ to $-1\leq\mu\leq 1$) to give $\widetilde{S}(\mu)=\sqrt{y}I(y)$, where $\widetilde{S}$ is degree $12k-$ 1 in $\mu$. One has from the condition $\sqrt{y}I(y)=\sqrt{y}+\mathcal{O}(y^{3/2})$ that $\widetilde{S}(\mu=0)=0,\widetilde{S}^{\prime}(\mu=0)=(2k-1)$. We compare this to the polynomial $T_{2k-1}(\mu)$, and see that one also has $T_{2k-1}(0)=0,T^{\prime}_{2k-1}(0)=(-1)^{k-1}(2k-1)$ (since $12k-$ 1 is odd) and thus $(-1)^{k-1}T_{2k-1}(\mu)-\widetilde{S}(\mu)$ possesses a double root at $\mu=0$. Then, since $\mu=0$ does not correspond to an extremum of $T_{2k-1}$, one may apply [4, Prop. 3.3] to show that $|\widetilde{S}(\mu)|>1$ on the interval $-1\leq\mu\leq 1$ unless it is equal to $(-1)^{k-1}T_{2k-1}(\mu)$ and thus the optimal polynomial $\sqrt{y}I(y)=(-1)^{k-1}T_{2k-1}(\sqrt{y}/(2k-1))$.
We may now proceed in a similar fashion to the $m$ odd case above, requiring now that $R(y)=0$ at the extrema of $T_{2k-1}$, which may be achieved via a polynomial of the form
$R(y)\propto\left(1-\frac{y}{(2k-1)^{2}}\right)\frac{d}{d\sqrt{y}}T_{2k-1}\left(\frac{\sqrt{y}}{2k-1}\right),$
where one may use the fact that $T^{\prime}_{2k-1}(x)=(2k-1)U_{2k-2}(x)$, and the condition $R(0)=1$ to determine the constant of proportionality, to give the optimal polynomial
$R(y)=(-1)^{k-1}\left(1-\frac{y}{(2k-1)^{2}}\right)U_{2k-2}\left(\frac{\sqrt{y}}{2k-1}\right),$ (10)
where $R(y)$ is indeed of degree $k$ in $y$. Taking $\sqrt{y}/(2k-1)\equiv\cos(\theta)$, one may then verify that $Q(y)=\cos^{2}((2k-1)\theta)+\sin^{2}(\theta)\sin^{2}((2k-1)\theta)\leq 1,0\leq y\leq(2k-1)^{2}$ as required. Consequently, one may obtain stability for $-(m-1)i\leq z\leq(m-1)i$ where $m=2k$ for the polynomial (replacing $z=i\sqrt{y}$) [9, Thm. 3]
$P(z)=(-1)^{k-1}\left[\left(1+\frac{z^{2}}{(2k-1)^{2}}\right)U_{2k-2}\left(\frac{z}{i(2k-1)}\right)+iT_{2k-1}\left(\frac{z}{i(2k-1)}\right)\right].$ (11)
It is then possible to assimilate eqs. 9 and 11 to the general form eq. 8. ∎
Remark 4.
In distinction to the case for theorems 1 and 2, integrators giving rise to stability polynomials are no longer optimal if one allows other competitor integrators; specifically, if one considers multi-step integrators, it is possible to reach $a=m$ [8, Sec. 5].
Remark 5.
The optimal $a=m-1$ result was first proven for $m$ odd (along with the construction of the polynomial eq. 9) in [13, Sec 4.3] and later for general $m$ in [14], with the two equivalent representations of the optimal polynomials for general $m$ finally given in [9, 12].
Note also that for $m$ odd, one has $\alpha=1/2$ as mentioned above, so that the integrator is in fact second order (although this is not the case for $m$ even, as a Taylor expansion of eq. 10 shows).
Acknowledgements
The author would like to thank Fernando Casas and Jesús María Sanz-Serna for helpful comments and discussions. LS has been supported by Ministerio de Ciencia e Innovación (Spain) through project PID2022-136585NB-C21, MCIN/AEI/10.13039/501100011033/FEDER, UE.